# A Review of Gravitational Memory and BMS Frame Fixing in Numerical Relativity

### v1

@@ VII.2 PN BMS frame
At the end of Sec. VII.1 the steps for mapping to the PN BMS frame through an iterative process using the BMS charges was outlined. Ultimately, to map to this frame one must simply find the BMS transformation which maps the charges of the NR system to match those of the PN system. This, however, requires that one has knowledge of the charges in PN theory. Unfortunately, most PN calculations focus on computing the strain and not the Weyl scalars, so the majority of the canonical BMS charges can not be computed in a PN formulation. Therefore, one can either compute these charges in PN or try to use alternative charges that are functions of only the strain. Because the later is easier, this is what is typically used in BMS frame fixing. But, even if one uses charges that are only functions of the strain, these charges must still be computed in PN. Fortunately, the center-of-mass charge, which is zero for PN systems, and the angular velocity vector, which can easily be computed numerically from the PN strain, are painless to obtain. The PN Moreschi supermomentum, however, because it involves a time integral, cannot be computed numerically and must instead be worked out analytically. This was first performed in Ref. Mitman et al. (2022), which computed it from the PN strain using Eq. (VII.1) to 3PN order without spins and 2PN order with spins. They then implemented this iterative procedure for fixing the frame of NR systems in the publicly available python module scri.
@@ VII.3 Superrest frame
Apart from comparing NR systems to PN predictions to construct more effective hybrid waveforms and models, NR simulations of black holes are also particularly useful for extracting the quasi-normal mode (QNM) amplitudes that are expected by perturbation theory. A subtle issue, however, is that when a perturbed black hole is formed, either through a black hole merger or stellar collapse, it is not in the frame that black hole perturbation theory is typically performed in. Put differently, the remnant black hole formed in some astrophysical event may not be described by the usual Kerr metric, but instead a boosted or a supertranslated version. Consequently, if one tries to study the perturbations of the remnant black hole formed in a NR simulation without accounting for the difference in BMS frames, then the analysis will fail in the same way that the hybridization shown in Fig. 6 fails.
Figure 8: Top row: The $(2,0)$ mode of the strain waveform for an NR system (black) and the best fit QNM model (blue). The QNM model contains seven $(2,0)$ overtones. On the left, the NR waveform is in the PN BMS frame, while on the right, it is instead in the superrest frame of the remnant black hole. Bottom row: Residual between the NR waveform and the fit.
In particular, in Fig. 8, we show exactly this. In the top row we show two fits of QNMs to the $(2,0)$ mode of the NR strain waveform from the simulation shown in Figs. 35, and 6. Meanwhile, in the bottom row we show the residuals. The QNM model that we consider has seven overtones and both mirror modes. [^22] We stress that here we use this model only for simplicity and to show the importance of fixing the frame. In reality, when trying to correctly fit an NR waveform with QNMs, one should take precaution to ensure that the QNMs are meaningfully resolved. As seen through the left-most column, which corresponds to the NR waveform in the PN BMS frame, i.e., the wrong frame for BHPT analyses, when the NR system is not in the superrest frame of the remnant black hole the strain can not be fit with QNMs, in part because of the nonzero asymptotic value resulting from the memory effect. But, if one maps the NR system to the superrest frame of the remnant black hole, then the QNM fit can be performed successfully and the residual is reduced by three orders of magnitude. This illustrates that fixing the frame is essential for extracting physical QNM amplitudes.
@@ VIII Discussion
In this review, we have presented a vast amount of information regarding memory effects and the BMS group. In particular, we began by providing some motivation for and physical intuition behind why one should study GWs at future null infinity, rather than at some point in the bulk of a NR simulation. Following this, we showed why the BMS group is the symmetry group of future null infinity and why it is important to understand the symmetries contained in the BMS group to ensure that GWs are modeled and studied accurately and robustly. Next, using this physical intuition for the BMS group that we developed, we provided some motivation for why memory effects can be expected to fall out of Einstein’s theory of GR due to conservation laws stemming from the BMS symmetries. This set the stage for our overarching literature review on memory effects and the BMS group as well as a more formal mathematical review of memory and the various BMS conservation laws.
After presenting this more introductory information, we then provided some discussion regarding memory effects and BMS frame fixing in the context of NR simulations. Specifically, we showed that the BMS conservations laws and the charges and fluxes that are associated with them provide a unique and helpful way to study a variety of features of NR waveforms. Furthermore, we showed that constraining the BMS freedoms of NR waveforms to match the canonical BMS frame of other waveform models, such as PN waveforms or black hole perturbation theory, is critically important for building accurate GW models as well as performing robust analyses with NR waveforms.
@@
It is our hope that this review will serve as a resource for those interested in learning more about memory effects and the BMS group in the context of NR simulations and GW modeling more generally. As new and more accurate GW detectors come online, understanding the complexities of GWs that GR predicts, such as those we have covered here, will be of the utmost importance, as they will likely be key to observing never-before-seen physics that will futher illuminate how our Universe works.
@@ Acknowledgments
The authors thank Eanna Flanagan for informative conversations about the history of memory effects and Alexander Grant for intuition behind BMS balance laws. Computations for this work were performed with the Wheeler cluster at Caltech. This work was supported by the Sherman Fairchild Foundation and NSF Grants No. PHY-2011968, PHY-2011961, PHY-2309211, PHY-2309231, OAC-2209656 at Caltech.
@@ A.0.1 $\mathrm{SL}(2,\mathbb{C})$ representation of a Lorentz transformation
Let us begin with rotations. For a rotation $R$ by an angle $\theta$ about the axis $\hat{r}=\left(r_{x},r_{y},r_{z}\right)$, one may write this as the quaternion
@@
Note that in Eq. (44) $\tilde{B}$ is a Hermitian matrix. Because $\tilde{L}$ is invertible, it has a unique polar decomposition. Correspondingly, a general Lorentz transformation can be written as a composition of a boost followed by a rotation, which in the $\mathrm{SL}(2,\mathbb{C})$ representation is

### v2

@@ VII.2 PN BMS frame
At the end of Sec. VII.1 the steps for mapping to the PN BMS frame through an iterative process using the BMS charges was outlined. Ultimately, to map to this frame one must simply find the BMS transformation which maps the charges of the NR system to match those of the PN system. This, however, requires that one has knowledge of the charges in PN theory. Unfortunately, most PN calculations focus on computing the strain and not the Weyl scalars, so the majority of the canonical BMS charges can not be computed in a PN formulation. Therefore, one can either compute these charges in PN or try to use alternative charges that are functions of only the strain. Because the later is easier, this is what is typically used in BMS frame fixing. But, even if one uses charges that are only functions of the strain, these charges must still be computed in PN. Fortunately, the center-of-mass charge, which is zero for PN systems, and the angular velocity vector, which can easily be computed numerically from the PN strain, are painless to obtain. The PN Moreschi supermomentum, however, because it involves a time integral, cannot be computed numerically and must instead be worked out analytically. This was first performed in Ref. Mitman et al. (2022), which computed it from the PN strain using Eq. (VII.1) to 3PN order without spins and 2PN order with spins. They then implemented this iterative procedure for fixing the frame of NR systems in the publicly available python module scri Boyle et al. (2024).
@@ VII.3 Superrest frame
Apart from comparing NR systems to PN predictions to construct more effective hybrid waveforms and models, NR simulations of black holes are also particularly useful for extracting the quasi-normal mode (QNM) amplitudes expected by perturbation theory Teukolsky (1973); Berti et al. (2009). A subtle issue, however, is that when a perturbed black hole is formed, either through a black hole merger or stellar collapse, it is not in the frame that black hole perturbation theory is typically performed in. Put differently, the remnant black hole formed in some astrophysical event may not be described by the usual Kerr metric, but instead a boosted or a supertranslated version. Consequently, if one tries to study the perturbations of the remnant black hole formed in a NR simulation without accounting for the difference in BMS frames, then the analysis will fail in the same way that the hybridization shown in Fig. 6 fails.
Figure 8: Top row: The $(2,0)$ mode of the strain waveform for a NR system (black) and the best fit QNM model (blue). The QNM model contains seven $(2,0)$ overtones. On the left, the NR waveform is in the PN BMS frame, while on the right, it is instead in the superrest frame of the remnant black hole. Bottom row: Residual between the NR waveform and the fit.
In particular, in Fig. 8, we show exactly this. In the top row we show two fits of QNMs to the $(2,0)$ mode of the NR strain waveform from the simulation shown in Figs. 35, and 6. Meanwhile, in the bottom row we show the residuals. The QNM model that we consider has seven overtones and both mirror modes. [^30] We stress that here we use this model only for simplicity and to show the importance of fixing the frame. In reality, when trying to correctly fit a NR waveform with QNMs, one should take precaution to ensure that the QNMs are meaningfully resolved. As seen through the left-most column, which corresponds to the NR waveform in the PN BMS frame, i.e., the wrong frame for BHPT analyses, when the NR system is not in the superrest frame of the remnant black hole the strain can not be fit with QNMs, in part because of the nonzero asymptotic value resulting from the memory effect. But, if one maps the NR system to the superrest frame of the remnant black hole, then the QNM fit can be performed successfully and the residual is reduced by three orders of magnitude. This illustrates that fixing the frame is essential for extracting physical QNM amplitudes.
@@ VIII Discussion
In this review, we have presented a vast amount of information regarding memory effects and the BMS group. In particular, we began in Sec. II.1 by studying where in a NR simulation one should extract gravitational wave information for gravitational wave detectors. While it may be tempting to extract GWs on some timelike wordline, like that of an inertial detector, we showed how this can quickly lead to coordinate ambiguities and thus motivated the need to extract GWs at future null infinity $\mathcal{I}^{+}$—the final destination of outgoing radiation. By working with data on $\mathcal{I}^{+}$, instead of in the bulk, one is free from the full diffeomorphism invariance of GR and obtains a more suitable set of coordinates to work with. However, as we outlined in Sec. II.2, even on $\mathcal{I}^{+}$ that are still troubling coordinate ambiguities that are present because of the symmetry group of null infinity: the BMS group. Consequently, to work with data on $\mathcal{I}^{+}$, one must have a robust understanding of the BMS group and the way in which it transforms coordinates and asymptotic data.
In Sec. II.3 we reviewed the infamous BMS group; specifically how it can be viewed as the Poincaré group, but with the usual spacetime translations replaced by an infinite group of transformations called supertranslations. These can be viewed as direction-dependent translations and arise from the fact that each point on $\mathcal{I}^{+}$, which can each be thought of as a single observer, ends up being casually disconnected from any other. Then, in Sec. II.5, we showed how an arbitrary BMS transformation changes the spacetime coordinates that the asymptotic data will be functions of. We also explained what this data on $\mathcal{I}^{+}$ should be, i.e., not only the gravitational wave strain, but also the five complex Weyl scalars, which encode information about the spacetime curvature.
Finally, in Sec. II.6, we used our understanding of the symmetries of $\mathcal{I}^{+}$ to construct an intuition for why memory effects should exist in GR. In particular, because of Neother’s theorem Wald and Zoupas (2000), one might expect that the supertranslation symmetries of $\mathcal{I}^{+}$ yield some kind of conservation law between an angle-dependent mass and an angle-dependent energy flux, like that of Eq. (II.6). However, because changes to the $\ell\geq 2$ harmonics of system’s mass multipole moment source GWs, one needs an extra ingredient on the left-hand side of Eq. (II.6), i.e., that of Eq. (II.6), to balance this fact. Therefore, the conservation law that stems from supertranslations, Eq. (II.6) or Eqs. (22), is really a statement about the memory that is induced on the spacetime by any system that emits gravitational radiation.
After our pedagogical overview, in Sec. III we reviewed the literature behind memory effects and the BMS group. Focusing on the memory first, we highlighted how this effect has been known since 1974, but did not receive much attention until its connection to asymptotic symmetries and soft theorems was found in 2013 and 2014 Strominger (2014); Winicour (2014); Flanagan and Nichols (2015); Strominger and Zhiboedov (2016). Ever since this realization, there has been a surge of interest and discovery in tangentially related fields, such as celestial holography Strominger (2017); Pasterski (2019); Raclariu (2021); Pasterski (2021); Pasterski et al. (2021).
In Sec. IV, we presented the more formal mathematics behind our earlier discussions. In particular, we showed how the Bondi-Sachs metric and Einstein’s equations yield a series of evolution equations for both the mass and angular momentum aspects (see Eqs. (18)), which, when integrated with respect to time, yield the formal version of the pedagogical conservation law of Eq. (II.6), i.e., Eqs. (22). These equations show that the real and imaginary parts of the gravitational wave strain (or shear) can be written in terms of mass and angular momentum charges and fluxes, which help illustrate the different ways in which the strain or memory can be sourced. That is, Eqs. (22) present the modern interpretation of memory: it can be sourced by a change in the system’s charges (linear or ordinary memory) or a change in the fluxes (nonlinear, Christodoulou, or null memory).
In Sec. VI, we implemented our mathematical results numerically for simulations of binary black hole mergers. In Sec. VI.1 we took Eqs. (22) and evaluated them for a GW150914-like NR simulation in Fig. 3. This showed how useful Eqs. (22) can be for studying NR simulations and extracting the memory contribution to the strain, as highlighted in the lower left panel of Fig. 3. In Sec. VI.2, in Fig. 4 we showed what the frequency spectrum of a typical memory signal looks like relative to the LIGO and CE amplitude spectral densities. As shown, even though memory is fairly pronounced in the time domain (see, e.g., Fig. 2), in the frequency domain memory occupies the lower part of the band which makes it challenging to detect in single events with current detectors. However, by stacking events and trying to detect memory in a population Hübner et al. (2020, 2021); Grant and Nichols (2023) or by relying on future detectors like CE, ET, or LISA Islo et al. (2019); Grant and Nichols (2023), the chances for observing memory for the first time are much higher.
Last, in Sec. VII we highlighted the importance of controlling the BMS freedoms for studying NR waveforms. First, in Fig. 6, we showed how NR and PN waveforms naturally disagree with one another, due to the fact that PN waveforms have information about the past history of the compact binary’s inspiral, while numerical simulations do not. Furthermore, Fig. 6 shows that, without resolving this issue, if one tries to construct a hybrid waveform which is needed for building waveform models for the LVK collaboration, then the hybrid waveform exhibits an unphysical feature due to this disagreement. Fortunately, this disagreement simply turns out to be an issue of the two waveforms being in different BMS frames, i.e., they differ by some arbitrary BMS transformation. In the text surrounding Fig. 6, we outlined how one can use BMS charges to determine this transformation and fix the BMS frame of the NR waveform to match that of the PN waveform. After doing so, one obtains Fig. 7 which shows much better agreement between NR and PN strain waveforms. Finally, in Fig. 8, we also highlighted why fixing the BMS frame is important for studying the ringdown phase of NR simulations with black hole perturbation theory. In particular, because of the memory, the remnant black hole formed in a black hole merger is not exactly Kerr, but supertranslated Kerr. Therefore, if one wishes to fit the ringdown phase with QNMs, they need to find the BMS transformation that maps the remnant black hole to the same frame of the usual Kerr metric: the superrest frame. After doing so, one can then perform meaningful ringdown fits and analyses, as is illustrated through the right-hand panels of Fig. 8.
@@
It is our hope that this review will serve as a resource for those interested in learning more about memory effects and the BMS group in the context of NR simulations and GW modeling more generally. As new and more accurate GW detectors come online, understanding the complexities of the GWs that GR predicts, such as those we have covered here, will be of the utmost importance, as they will likely be key to observing never-before-seen physics that will futher illuminate how our Universe works.
@@ Acknowledgments
The authors thank Eanna Flanagan for informative conversations about the history of memory effects and Alexander Grant for intuition behind BMS balance laws. Computations for this work were performed with the Wheeler cluster at Caltech. This work was supported by the Sherman Fairchild Foundation and NSF Grants No. PHY-2011968, PHY-2011961, PHY-2309211, PHY-2309231, OAC-2209656 at Caltech. The work of L.C.S. was partially supported by NSF CAREER Award PHY-2047382 and a Sloan Foundation Research Fellowship.
@@ A.0.1 $\mathrm{SL}(2,\mathbb{C})$ representation of a Lorentz transformation
Let us begin with rotations. For a rotation $R$ by an angle $\theta$ about the axis $\hat{r}=\left(r_{x},r_{y},r_{z}\right)$, one may write this as the quaternion [^31] For a review of quaternions, see Ref. Doran and Lasenby (2007).
@@
Note that in Eq. (46) $\tilde{B}$ is a Hermitian matrix. Because $\tilde{L}$—the $\mathrm{SL}(2,\mathbb{C})$ representation of some Lorentz transformation—is invertible, it has a unique polar decomposition. Correspondingly, a general Lorentz transformation can be written as a composition of a boost followed by a rotation, which in the $\mathrm{SL}(2,\mathbb{C})$ representation is